Samples of weights of 200 students each are drawn from a normal population of the weights of students with S.D. 10

pound. Variance in each sample is computed. Find (i) mean (ii) standard deviation of the sampling distribution of sample

variance.


1
Expert's answer
2021-12-29T17:18:34-0500

From the central limit theorem we see that the sampling distribution of the mean approaches a normal distribution with a mean of


"E(\bar{X})=\mu_{\bar{X}}=\mu=50 \ kg"

and a variance of



"{\sigma^2 \over n}={5^2 \over 30}={5 \over 6}"

The standard deviation of the sampling distribution



"\sigma_{\bar{X}}={5 \over \sqrt{30}}={\sqrt{30}\over 6}"